# MindShift: Leveraging Large Language Models for Mental-States-Based Problematic Smartphone Use Intervention

### v1

@@ Abstract.
Problematic smartphone use negatively affects physical and mental health. Despite the wide range of prior research, existing persuasive techniques are not flexible enough to provide dynamic persuasion content based on users’ physical contexts and mental states. We first conduct a Wizard-of-Oz study (N=12) and an interview study (N=10) to summarize the mental states behind problematic smartphone use: boredom, stress, and inertia. This informs our design of four persuasion strategies: understanding, comforting, evoking, and scaffolding habits. We leverage large language models (LLMs) to enable the automatic and dynamic generation of effective persuasion content. We develop MindShift, a novel LLM-powered problematic smartphone use intervention technique. MindShift takes users’ in-the-moment physical contexts, mental states, app usage behaviors, users’ goals & habits as input, and generates high-quality and flexible persuasive content with appropriate persuasion strategies. We conduct a 5-week field experiment (N=25) to compare MindShift with baseline techniques. The results show that MindShift significantly improves intervention acceptance rates by 17.8-22.5% and reduces smartphone use frequency by 12.1-14.4%. Moreover, users have a significant drop in smartphone addiction scale scores and a rise in self-efficacy. Our study sheds light on the potential of leveraging LLMs for context-aware persuasion in other behavior change domains.
@@
ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
In recent years, the ubiquitous presence of smartphones has increased people’s reliance on digital devices. This trend frequently leads to problematic smartphone usage behaviors (Hwang et al., 2012), especially among adolescents and young adults (Lee et al., 2014). Prior studies suggest that problematic smartphone usage can detrimentally affect people in various areas such as efficiency (leading to diminished academic or work performance (Arefin et al., 2018; Duke and Montag, 2017)), physical well-being (resulting in decreased sleep and activity levels (Clayton et al., 2015; Lapointe et al., 2013; Xu et al., 2019, 2021)), and mental health (manifesting as anxiety and depression (Hartanto and Yang, 2016; Harwood et al., 2014; Xu et al., 2022a)). Many individuals have recognized their problematic smartphone usage and sought to reduce their over-reliance on smartphones (Hiniker et al., 2016a; Ko et al., 2015).
There has been a plethora of academic research and commercial products that provide just-in-time (JIT) smartphone use interventions that intervene at the moment when problematic use is taking place (Nahum-Shani et al., 2018). Existing intervention techniques can be summarized into four categories based on their degree of enforcement: (1) Feedback: offering insights on phone usage patterns, enhancing user awareness about their smartphone habits (Whittaker et al., 2016; Kim et al., 2016; Collins et al., 2014; Andone et al., 2016); (2) Reminders: aiming to counter immediate phone indulgence and promote self-reflection through pop-up notifications (Hiniker et al., 2016a; Park and Gweon, 2015); (3) Interaction friction: raising the effort needed to use the phone, thereby reducing its allure by introducing tasks like typing (Park et al., 2018; Okeke et al., 2018; Xu et al., 2022b); (4) Lockout: disabling the user’s phone access for a specified duration(Kim et al., 2019a; Ko et al., 2015; Löchtefeld et al., 2013; Kim et al., 2019b).
@@
However, we notice a few gaps among existing intervention techniques. One major issue is the specific moment of intervention. Most existing intervention techniques are time-based, such as simply upon app launch (Xu et al., 2022b; Kim et al., 2019b), or after hitting the upper limit of usage duration (Hiniker et al., 2016a; Kim et al., 2019a; Park et al., 2018). However, these methods ignore an important factor: users’ smartphone use purpose. Users’ smartphone use purposes can be generally divided into instrumental use (i.e., for specific goals such as working or intentional relaxation) and habitual use (i.e., ritualistic behavior, without a clear goal, such as passive social media content consumption) (Rubin, 2009; Hiniker et al., 2016b). Recent research suggests intervention should mainly focus on habitual use since it often gives people a lower sense of meaningfulness (Lukoff et al., 2018). Thus, an intervention technique should consider not only the use time but also the smartphone usage purpose.
Furthermore, we identify a further gap in the intervention content. Existing JIT intervention techniques rarely take users’ in-the-moment contexts (e.g., current environment, mental states) into account to achieve context-aware intelligent intervention (Thomas Craig et al., 2021). This is particularly important for smartphone persuasion intervention, one of the reminder-based techniques that show messages designed to convince users to stop using smartphones. The persuasion content is the key to the effectiveness of intervention. Although some intervention techniques are designed to be context-aware (e.g., an intervention is introduced when users are around a specific location (Kim et al., 2017; Ko et al., 2015)), they are simply rule-based and cannot adapt to the dynamic changing physical environment. In addition, no prior work considers users’ mental state, which is another important aspect of user contexts.
@@
To address these gaps, we first conducted a Wizard-of-Oz study (N=12), followed by an interview study (N=10) to better understand users’ mental states during problematic phone usage. Focusing on habitual usage, we summarized three major mental states to address: boredom, stress, and inertia. Leveraging the Dual Systems Theory (Hofmann et al., 2009) and the ERG (Existence, Relatedness, and Growth) Theory (Caulton, 2012), we proposed a set of four persuasion strategies and built the mapping between these strategies to the three mental states. These strategies include 1) understanding, 2) comforting, 3) evoking, and 4) scaffolding habits.
@@
Based on our persuasion strategies, we designed and implemented MindShift (Figure 1), a new JIT intervention technique that can provide dynamic, personalized persuasion content based on user contexts. MindShift leverages Large Language Models’ (LLMs) strong capability in commonsense comprehension and natural language generation (Chowdhery et al., 2022; Min et al., 2021). Specifically, we employ LLMs to generate proper and effective persuasive content based on phone usage behavior, physical contexts, mental states, goals and habits with appropriate persuasion strategies we design.
@@
To evaluate the effectiveness of MindShift, we conducted a 5-week field experiment to deploy our intervention to 25 participants. We compared MindShift against the basic notification-based intervention that asked users to reflect and report their smartphone usage purpose. Moreover, to better evaluate the effect of the mental states factor, we also compared MindShift against a simplified version that excludes the mental states factor from the LLM-based content generation.
@@
Our study results indicate that MindShift significantly outperforms the baseline method by 17.8-22.5% on the intervention acceptance rate. MindShift also reduces overall app opening frequency by 12.1-14.4% and app usage duration by 2.4-9.8%. Specifically, MindShift leads to a significant reduction in habitual proportions by 6.8%. Comparing the two versions of MindShift, including the mental states factor in the persuasion content generation can significantly enhance the acceptance rate by 8.1%. The subjective report data show that participants have a significant decrease (25.8-34.7%) in the smartphone addiction scale (SAS) score and a significant increase (10.4-10.7%) in the self-efficacy scale score after using MindShift. Participants also expressed a strong willingness to continue using this effective intervention approach.
The contribution of our paper can be summarized as follows:
@@
    We conducted a Wizard-of-Oz study and an interview study to investigate the mental states that end-users experience during smartphone habitual use. Our study highlights three major mental states (boredom, stress, and inertia), leading us to design four persuasion strategies grounded in the Dual Systems Theory and the ERG Theory: Understanding, Comforting, Evoking, and Scaffolding Habits.
@@
    We created MindShift, a novel persuasive intervention technique that leveraged an LLM to generate dynamic and personalized persuasion content based on appropriate persuasion strategies, users’ phone usage behavior, physical contexts, mental states, goals and habits.
@@
    We conducted a field experiment by deploying MindShift. Our results indicate that our new technique could significantly improve intervention acceptance rates and reduce smartphone use. Users’ subjective feedback also corroborates these observations, demonstrating the effectiveness of MindShift.
@@ 2. Related work
In this section, we first introduce the definition of smartphone problematic use and habitual phone use (Sec. 2.1). We then briefly overview existing intervention techniques (Sec. 2.2). Close to our work, we introduce the behavior change persuasion techniques and their relationship with the boost of LLMs (Sec. 2.3). Finally, we cover the theoretical foundations and mental models of problematic smartphone use (Sec. 2.4).
2.1. Smartphone Problematic Use
Many studies have explored the definition of problematic smartphone use, which can be broadly classified into two categories. The first category defines whether users exhibit addictive behaviors toward their phones. Some studies assessed addictive behavior by measuring the level of user dependence on smartphones through questionnaires, such as the Smartphone Addiction Scale (SAS) and the Smartphone Addiction Inventory (SPAI) (Lin et al., 2014; Kwon et al., 2013). The second category defines whether a specific instance of phone use is problematic. Recent research suggests that problematic smartphone use is determined not only by excessive use but also by the purpose and content of use in specific situations (Harwood et al., 2014; Prasad et al., 2021). Studies have pointed out that users’ purposes for phone use can be categorized into (1) habitual use that is performed unconsciously and ritualistically usually without a specific goal (Pinder et al., 2018), and (2) instrumental use with a specific task or goal in mind (Lukoff et al., 2018). Existing research suggests that habitual use should be the primary target for intervention (Morris et al., 2018; Lukoff et al., 2018; Pinder et al., 2018). In this paper, we combine these two definitions. We use SAS to measure the level of addictive smartphone usage behavior. We also distinguish users’ phone use intention and focus on intervening habitual use.
@@ 2.2. Problematic Smartphone Use Intervention
Existing problematic smartphone use intervention techniques can be roughly divided into four categories based on their level of enforcement: The first category provides users with information about their behavior (Whittaker et al., 2016; Kim et al., 2016; Hiniker et al., 2016a; Collins et al., 2014; Apple, 2021; Lyngs et al., 2022; Nwagu, 2023), requiring users to view it themselves to increase awareness of phone usage. The second actively sends notifications and reminders to users to provoke their reflection such as reminding users of their daily goals (Hiniker et al., 2016a), sending them the results of previous decision-balancing exercises (Park and Gweon, 2015), or informing their usage time (Apple, 2021). This category presents text to users and, therefore, serves as a persuasion. The third involves increasing the difficulty of using the phone and intentionally slowing down user interactions to suppress the desire to use it. This can include requiring users to enter random numbers or engage in self-reflection (Park et al., 2018; Xu et al., 2022b), keeping the phone vibrating continuously (Okeke et al., 2018). The fourth is particularly forceful by directly locking the users’ apps or phone for a specific duration (APP, 2021; Kim et al., 2019a; Ko et al., 2015).
There are concerns about these methods’ ability to strike an optimal balance between user experience and effectiveness. The first two intervention categories (i.e., feedback-based self-awareness and notification reminders) rely on the user’s self-control and can be easily ignored, leading to limited effectiveness (Whittaker et al., 2016; Kim et al., 2016; Hiniker et al., 2016a). While the other two intervention types (i.e., increasing usage difficulty and locking out) are relatively more effective, research has shown that users often struggle to maintain compliance due to a diminished user experience (Kim et al., 2019a). Once the restrictions are removed, users often relapse and revert to problematic use behavior in a short period of time (Kim et al., 2019a; Xu et al., 2022b). To bridge the gap, we hope to develop a method that ensures a good user experience while boosting intervention effectiveness.

### v2

@@ Abstract.
Problematic smartphone use negatively affects physical and mental health. Despite the wide range of prior research, existing persuasive techniques are not flexible enough to provide dynamic persuasion content based on users’ physical contexts and mental states. We first conducted a Wizard-of-Oz study (N=12) and an interview study (N=10) to summarize the mental states behind problematic smartphone use: boredom, stress, and inertia. This informs our design of four persuasion strategies: understanding, comforting, evoking, and scaffolding habits. We leveraged large language models (LLMs) to enable the automatic and dynamic generation of effective persuasion content. We developed MindShift, a novel LLM-powered problematic smartphone use intervention technique. MindShift takes users’ in-the-moment app usage behaviors, physical contexts, mental states, goals & habits as input, and generates personalized and dynamic persuasive content with appropriate persuasion strategies. We conducted a 5-week field experiment (N=25) to compare MindShift with its simplified version (remove mental states) and baseline techniques (fixed reminder). The results show that MindShift improves intervention acceptance rates by 4.7-22.5% and reduces smartphone usage duration by 7.4-9.8%. Moreover, users have a significant drop in smartphone addiction scale scores and a rise in self-efficacy scale scores. Our study sheds light on the potential of leveraging LLMs for context-aware persuasion in other behavior change domains.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the CHI Conference on Human Factors in Computing Systems; May 11–16, 2024; Honolulu, HI, USAbooktitle: Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24), May 11–16, 2024, Honolulu, HI, USAdoi: 10.1145/3613904.3642790isbn: 979-8-4007-0330-0/24/05ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
(Left) A picture illustrating that a user is problematically using her phone. (Medium) A flowchart illustrating the MindShift system. The left side of the flowchart is the system’s inputs, organized vertically into two main sections. The top section, labeled ’Real-time information,’ details two factors: (1) ’Usage Behavior’ (containing frequency and duration of phone usage), and (2)’Physical Context’ (describing the time and location of the user). Below this, the second section is titled ’Design from Formative Study.’ It includes three variables: (1)’Mental State’ (options are Stress, Boredom, or Inertia), (2)’Persuasion Strategy’ (options are Understanding, Comforting, Evoking, or Scaffolding Habits), and (3)’Goals and habits.’ An arrow links these inputs to an icon representing a ’Large Language Model’. On the right side of the flowchart, a depiction shows a user receiving a personalized message from the system. (Right) A picture illustrating that the user stops playing with her phone.
In recent years, the ubiquitous presence of smartphones has increased people’s reliance on digital devices, resulting in problematic smartphone usage behaviors, i.e., excessive or mindless usage with negative consequences (Lukoff et al., 2018; Hwang et al., 2012), especially among adolescents and young adults (Lee et al., 2014). Prior studies suggest that problematic smartphone usage can detrimentally affect people in various areas such as efficiency (leading to diminished academic or work performance (Arefin et al., 2018; Duke and Montag, 2017)), physical well-being (resulting in decreased sleep and activity levels (Clayton et al., 2015; Lapointe et al., 2013; Xu et al., 2019, 2021)), and mental health (manifesting as anxiety and depression (Hartanto and Yang, 2016; Harwood et al., 2014; Xu et al., 2022a)). Many individuals have recognized their problematic smartphone usage and sought to reduce their over-reliance on smartphones (Hiniker et al., 2016a; Ko et al., 2015).
A plethora of academic research and commercial products provide just-in-time (JIT) smartphone use interventions, intervening precisely when problematic use occurs (Nahum-Shani et al., 2018). These interventions fall into four categories based on their enforcement levels: (1) Self-monitoring: offering insights on phone usage patterns via notification or visualization, enhancing user awareness about their smartphone habits (Whittaker et al., 2016; Kim et al., 2016; Collins et al., 2014; Andone et al., 2016); (2) Reminders: countering immediate phone indulgence and promoting self-reflection through pop-up notifications (Hiniker et al., 2016a; Park and Gweon, 2015); (3) Interaction friction: raising the effort needed to use the phone, thereby reducing its allure by introducing tasks like typing (Park et al., 2018; Okeke et al., 2018; Xu et al., 2022b); (4) Lockout: disabling the user’s phone access for a specified duration (Kim et al., 2019a; Ko et al., 2015; Löchtefeld et al., 2013; Kim et al., 2019b). However, there are a few gaps among existing intervention techniques.
First, existing methods are limited to strike a balance between effective intervention engagement and good usability (Monge Roffarello and De Russis, 2019). The first two categories rely on the user’s self-control and are easily ignored, leading to low engagement and limited effectiveness (Whittaker et al., 2016; Kim et al., 2016; Hiniker et al., 2016a). The other two types are more restrictive, often causing user frustration due to reduced usability across different contexts (Kim et al., 2019a). Our approach utilizes reminder-based interventions with persuasive content to encourage reduced smartphone usage. Persuasion, typically through natural language to influence people’s thoughts and behavior (Simons, 1976; Rapp, 2002; Cialdini and Cialdini, 2007), is more effective with diverse and context-specific content (Kaptein et al., 2015, 2012), as supported by recent studies highlighting the success of personalized (Kovacs et al., 2018; Choe et al., 2017; Harrison et al., 2015; Kaptein et al., 2012), context-aware (Kim et al., 2017; Ko et al., 2015; Thomas Craig et al., 2021; Purohit and Holzer, 2019) interventions. However, most current reminders use repetitive, template-based content, reducing efficacy (Apple, 2021; Inc, 2023; Xu et al., 2022b). To overcome this, we employ Large Language Models (LLMs) (OpenAI, 2022; Chowdhery et al., 2022) to generate varied persuasive content. LLMs’ reasoning ability provides a promising solution to infer users’ current activities based on contextual information collected from smartphone sensors, such as time and location (Chen et al., 2023; Kim et al., 2024; Englhardt et al., 2023), enabling the creation of more relevant and effective intervention language.
Second, we identified the opportunity to leverage mental states associated with problematic smartphone use, an essential aspect of user contexts. While some current interventions use context-based strategies, like triggering interventions at specific times and locations (Kim et al., 2017; Ko et al., 2015), they tend to focus mainly on external physical contexts, neglecting internal mental factors. Mental factors like stress and negative emotions are increasingly recognized as one key factor leading to problematic smartphone use  (Lyngs et al., 2019; Duvenage et al., 2020; Wang et al., 2014; Wang and Lee, 2020). However, existing studies primarily address prolonged mental states rather than momentary contexts. Our study aims to bridge this gap by integrating an understanding of in-the-moment mental states into the intervention framework. We believe that a more holistic approach, considering both the physical and mental contexts, could enhance intervention effectiveness.
@@
To address these gaps, we first conducted a Wizard-of-Oz study (N=12), followed by an interview study (N=10) to better understand users’ mental states during problematic phone usage. Focusing on habitual usage (i.e., ritualistic behavior, without a clear goal, such as passive social media content consumption) (Rubin, 2009; Hiniker et al., 2016b), we summarized three major mental states to address: boredom, stress, and inertia. Building on the Dual Systems Theory (Hofmann et al., 2009) and the ERG (Existence, Relatedness, and Growth) Theory (Caulton, 2012), we proposed four persuasion strategies: 1) understanding, 2) comforting, 3) evoking, and 4) scaffolding habits.
@@
Integrating our persuasion strategies with LLMs, we designed and implemented MindShift (Figure 1), a new JIT intervention technique that can provide dynamic, personalized persuasion content based on user contexts. MindShift leverages LLMs’ strong capability in commonsense comprehension and natural language generation (Chowdhery et al., 2022; Min et al., 2021) to generate proper and effective persuasive content based on real-time information (phone usage behavior, physical contexts, and mental states) and long-term user states (user goals and habits), guided by persuasion strategies we designed.
@@
To evaluate MindShift’s effectiveness, we conducted a 5-week field experiment deploying our intervention to 25 participants. We compared MindShift against the baseline, a basic notification-based intervention that asked users to reflect on and report the purpose of their smartphone usage. Moreover, to assess the effect of the mental states factor, we compared MindShift against a simplified version, MindShift-Simple, that excludes the mental states factor from the LLM-based content generation.
@@
Our study results indicate that MindShift and MindShift-Simple outperformed the baseline method on the intervention acceptance rate by 22.5% and 17.8%, respectively, with statistical significance. They also significantly reduce overall app opening frequency by 12.1% and 14.4% and app usage duration by 9.8% and 2.4%. Comparing MindShift and MindShift-Simple, including the mental state factor enhances the persuasion acceptance rate by 8.1% with statistical significance. Moreover, the subjective report data shows that participants using MindShift and MindShift-Simple experience a significant reduction in smartphone addiction scale (SAS) score (34.7% and 25.8% respectively) and an increase in the self-efficacy scale score (10.7% and 10.4% respectively).
Our paper makes the following contributions:
@@
    We conducted a Wizard-of-Oz study and an interview study, uncovering three major mental states (boredom, stress, and inertia) during habitual smartphone use, which led us to design four persuasion strategies grounded in the Dual Systems Theory and the ERG Theory: Understanding, Comforting, Evoking, and Scaffolding Habits.
@@
    We created MindShift, a novel persuasive intervention technique leveraging LLMs to generate dynamic and personalized persuasion content based on users’ phone usage behavior, physical contexts, mental states, goals and habits, and appropriate persuasion strategies.
@@
    We conducted a field experiment by deploying MindShift, demonstrating significant improvements in intervention acceptance rates and reduced smartphone use by Mindshift. Users’ subjective feedback also corroborated these observations, validating the effectiveness of MindShift.
@@ 2. Related work
In this section, we define problematic smartphone use and habitual smartphone use (Sec. 2.1), explore the reasons behind engagement in problematic smartphone use (Sec. 2.2). We then briefly overview existing intervention techniques (Sec. 2.3). Finally, close to our work, we introduce behavior change persuasion techniques and their relationship with the emergence of LLMs (Sec. 2.4).
2.1. Problematic Smartphone Use
Many studies have explored the definition of problematic smartphone use, which can be broadly classified into two categories. The first category defines whether users exhibit addictive behaviors toward their phones. Some studies assess addictive behavior by measuring the level of user dependence on smartphones through questionnaires, such as the Smartphone Addiction Scale (SAS) and the Smartphone Addiction Inventory (SPAI) (Lin et al., 2014; Kwon et al., 2013). The second category defines whether a specific instance of phone use is problematic. Growing research suggests that problematic smartphone use is determined not only by excessive use but also by the purpose and content of use in specific situations (Harwood et al., 2014; Prasad et al., 2021; Purohit et al., 2020; Lukoff et al., 2023, 2021, 2018). Studies have indicated that phone use purposes can be categorized into (1) habitual use that is performed unconsciously and ritualistically usually without a specific goal (Pinder et al., 2018), and (2) instrumental use with a specific task or goal in mind (Lukoff et al., 2018). Existing research suggests that habitual use should be the primary target for intervention (Morris et al., 2018; Lukoff et al., 2018; Pinder et al., 2018). In this paper, we use SAS to measure users’ level of addictive smartphone usage. We also distinguish users’ phone use purpose and focus on intervening habitual use.
2.2. Understanding Problematic Smartphone Use through A Dual Systems Perspective
Understanding what leads to problematic smartphone use is essential for the design of effective persuasion strategies. The Dual Systems Theory (Hofmann et al., 2009; Kahneman, 2011) has been used to explain the phone usage patterns (Pinder et al., 2018). This theory divides human cognitive activities into two types: System 1 (fast, intuitive, unconscious) and System 2 (slow, analytical, conscious). Problematic smartphone use is typically driven by System 1 (Hagger, 2016), as it mainly involves unconscious, rapid responses and is easy to be guided by instant gratification. Research suggests that two key factors contribute to the failure of users to act on their goals: (1) limited ability of System 2 control; and (2) fluctuations of System 2 caused by emotional states and fatigue (Lyngs et al., 2019). For the first factor, a growing amount of research suggests that the limited ability of control is attributed more to apps’ deliberate design than to users themselves (Purohit et al., 2023b; Baughan et al., 2022; Lukoff et al., 2021; Monge Roffarello et al., 2023; Gray et al., 2018). Recent work identifies types of attention-capture deceptive designs in digital interfaces, such as neverending autoplay and infinite scroll (Monge Roffarello et al., 2023). For the second factor, some findings suggest that mental states play a significant role in habitual smartphone use (Pinder et al., 2018; Wang et al., 2014; Wang and Lee, 2020) and previous research has identified external contexts, such as social awkwardness (Tran et al., 2019; Purohit et al., 2020), that may trigger the habitual use. However, there is limited research exploring what specific kinds of users’ in-the-moment mental states behind habitual use. In our work, we pinpoint the major mental states linked to habitual use and propose corresponding persuasion strategies.
@@ 2.3. Problematic Smartphone Use Intervention
Existing problematic smartphone use intervention techniques fall into two groups: external interventions that monitor and limit use, and internal interventions that change the interface itself.
External intervention can be roughly divided into four categories based on enforcement level: The first category provides users with information about their behavior such as visualization of usage (Whittaker et al., 2016; Kim et al., 2016; Hiniker et al., 2016a; Collins et al., 2014; Apple, 2021; Lyngs et al., 2022; Nwagu, 2023; Orzikulova et al., 2024; Lu et al., 2024), requiring users to view it themselves to increase awareness of phone usage. The second actively sends reminders to users to provoke their reflection such as reminding users of their daily goals (Hiniker et al., 2016a; Lyngs et al., 2020), informing their usage time (Apple, 2021) or the number of opens (Purohit et al., 2023a). This category presents text to users and, therefore, serves as a persuasion. The third involves increasing the difficulty of using the phone and intentionally slowing down user interactions to suppress the desire to use it, such as requiring users to enter random numbers or type self-reflective text (Park et al., 2018; Xu et al., 2022b) and keeping the phone vibrating continuously (Okeke et al., 2018). The fourth is particularly forceful by directly locking the users’ apps or phones for a specific duration (APP, 2021; Kim et al., 2019a; Ko et al., 2015). There are concerns about these methods’ ability to strike an optimal balance between usability and effectiveness (Monge Roffarello and De Russis, 2019).
Internal intervention involves redesigning app interfaces to counteract attention-capturing deceptive designs (Monge Roffarello et al., 2023). For example, increasing user awareness of time spent through reading history labels  (Baughan et al., 2022) and specific color change  (Monge Roffarello and De Russis, 2023), eliminating the addictive design of infinite scroll through removing  (Lyngs et al., 2020) and adjusting the newsfeed  (Kim et al., 2016; Purohit et al., 2023b; Zhang et al., 2022), decreasing the guilty pleasure recommendations through using adaptable commitment interface  (Lukoff et al., 2023) and redesigning search interface  (Monge Roffarello and De Russis, 2023). Compared to external intervention, internal intervention can better balance effectiveness and long-term experience (Zhang et al., 2022; Purohit et al., 2023b). However, these internal methods often require third-party development, as large companies rarely adopt such designs themselves due to financial interests (Gray et al., 2018). This necessitates additional development costs and the proposed design is typically tailored for a single app, making it hard to apply broadly.
Therefore, we hope to create a universal external intervention, using the form of reminders to ensure usability while boosting intervention effectiveness through personalized persuasion.